# Conformal Prediction in Hierarchical Classification

### v1

@@ 3 Proposed methods
Building upon the concepts in Section 2.3, we are now ready to discuss two algorithms that provide valid set-valued predictions for hierarchical classification that satisfy (1) above. The first algorithm provides marginal validity guarantees in classical hierarchical classification settings, i.e. where sets are restricted to nodes of the tree structure, thus $R_{\mathcal{T}}(\hat{Y})=1$. The second algorithm provides marginal validity guarantees in restricted set-valued prediction settings, i.e., where the representation complexity of the predicted set is restricted to $R_{\mathcal{T}}(\hat{Y})\leq r$ for a user-defined value $r$. Note that traditional conformal prediction in flat classification is obtained for the second algorithm when there is no restriction on the representation complexity (e.g. $r=K$).
3.1 Conformal restricted set-valued prediction (CRSVP)
The first approach, called conformal restricted set-valued prediction (CRSVP), predicts sets that are restricted to nodes of the hierarchy, thus having representation complexity $R_{\mathcal{T}}(\hat{Y})=1$. Assume one has fitted a hierarchical classifier using a training dataset, as described in Section 2.1 and let $\hat{y}(\bm{x})$ denote the mode of the distribution $\hat{P}(.\,|\,\bm{x})$, i.e. the leaf node with highest probability mass. For our first approach, we consider the following set-valued predictor:
$\hat{Y}_{1}(\bm{x},u,\tau)=\operatorname*{\arg\max}_{\hat{Y}\in\mathrm{Path}(\hat{y}(\bm{x}))}\big{\{}|\hat{Y}|:\hat{P}(\hat{Y}\,|\,\bm{x})+$
$u\cdot\hat{P}(\mathrm{pa}(\hat{Y})\setminus\hat{Y}\,|\,\bm{x})\leq\tau\big{\}}\,,$ (7)
with $\mathrm{Path}(v)$ the path to the root, and $\mathrm{pa}(v)$ the parent of a node $v$, as defined in Section 2. $\hat{P}(\hat{Y}\,|\,\bm{x})$ can be computed using (2), but also observe that the probability mass of an internal node $v$ corresponds to $\hat{P}(v\,|\,\bm{x})=\sum_{c\in v}\hat{P}(c\,|\,\bm{x})$, i.e., the sum of the probability masses of the leaf nodes that are descendants of $v$. However, computing the probability mass of an internal node and its associated set of classes in this way is computationally less efficient than using the chain rule in (2), when hierarchical classifiers have been fitted during training.
It is clear that the set-valued predictor in (3.1) satisfies the nested property in (4). Indeed, starting from the mode of the distribution $\hat{y}(\bm{x})$, every node on the path connecting that node and the root is nested by definition of the hierarchical tree structure $\mathcal{T}$. In line with Angelopoulos et al. (2020), the first term increases as we move up the tree, while the second term contains a random draw from the uniform distribution $\mathcal{U}(0,1)$, for handling discrete jumps in probability mass when following the path towards the root. The randomization is needed to prevent over-coverage.
The algorithm for calculating the threshold in (2.3) using the calibration dataset is presented in Algorithm 1. This algorithm first computes the mode of the estimated class probabilities for each instance in the calibration dataset. For small hierarchies, this can be done using exhaustive search, but more efficient inference algorithms have been developed when the computational cost becomes a burden (Dembczyński et al., 2012; Kumar et al., 2013; Mena Waldo et al., 2015). Subsequently, starting from the mode, Algorithm 1 computes for every calibration instance the internal node, on the path to the root, that also includes the true class label. Nonconformity scores that incorporate a randomization component are constructed for these internal nodes, and the final threshold is deduced from the scores. Algorithm 1 has a worst case $\mathcal{O}(N\,K)$ time complexity when exhaustive search is applied in the first step.
The pseudocode for computing the set-valued prediction for a new test instance, as defined in (3.1), is shown in Algorithm 2. Given a hierarchical classifier and a balanced hierarchy, the computational complexity during test time is given by $\mathcal{O}(\log K)$.
Algorithm 1 CRSVP calibration – Input: $\{(\bm{x}_{i},y_{i},u_{i})\}_{i=1}^{N},\hat{P},\mathcal{V}_{\mathcal{T}}$, Output: Threshold in (2.3).
@@ 3.1 Conformal restricted set-valued prediction (CRSVP)
5:       $\hat{p}_{\hat{Y}^{\prime}}\leftarrow\hat{p}_{\hat{Y}}$
6:       $\hat{Y}\leftarrow\mathrm{pa}(\hat{Y}),\hat{p}_{\hat{Y}}\leftarrow\hat{P}(\hat{Y}\,|\,\bm{x}_{i})$
@@
10:  $\tau^{*}\leftarrow$ the $\lceil(1-\alpha)(N+1)\rceil$-th largest value in $\{\tau_{i}\}_{i=1}^{N}$
@@
Algorithm 2 CRSVP inference – Input: $\bm{x},\tau^{*},u,\hat{P},\mathcal{V}_{\mathcal{T}}$, Output: Set-valued prediction in (3.1).
1:  $\hat{Y}\leftarrow\operatorname*{\arg\max}_{c\in\mathcal{Y}}\hat{P}(c\,|\,\bm{x}),\hat{Y}^{\prime}\leftarrow\emptyset$
2:  $\hat{p}_{\hat{Y}}\leftarrow\hat{P}(\hat{Y}\,|\,\bm{x}),\hat{p}_{\hat{Y}^{\prime}}\leftarrow 0$
3:  while $|\hat{Y}^{\prime}|\neq K$ do
4:    if $\hat{p}_{\hat{Y}}+u\cdot\hat{P}(\mathrm{pa}(\hat{Y})\setminus\hat{Y}\,|\,\bm{x})>\tau^{*}$ then
@@
7:    $\hat{Y}^{\prime}\leftarrow\hat{Y},\hat{Y}\leftarrow\mathrm{pa}(\hat{Y})$
8:    $\hat{p}_{\hat{Y}^{\prime}}\leftarrow\hat{p}_{\hat{Y}},\hat{p}_{\hat{Y}}\leftarrow\hat{P}(\hat{Y}\,|\,\bm{x})$
@@
10:  return $\hat{Y}^{\prime}$
3.2 Conformal set-valued prediction with representation complexity (CRSVP-$r$)
It should be obvious for the reader that sets of representation complexity one quickly become very big, since they correspond to internal nodes of the hierarchy that exhibit a predefined coverage level. Therefore, we present a second approach, called conformal set-valued prediction with representation complexity (CRSVP-$r$), that relaxes the representation complexity constraint to $R_{\mathcal{T}}(\hat{Y})\leq r$ with $r$ a user-defined parameter. Let
$\hat{\omega}(y;\bm{x})=\left\{y^{\prime}\in\mathcal{Y}:\hat{P}(y^{\prime}\,|\,\bm{x})\geq\hat{P}(y\,|\,\bm{x})\right\}$ (8)
be the set of classes that are more likely than the label $y$ for $\bm{x}$. As a crucial concept for the second approach, we first define the following optimization problem:
$A_{r}(y;\bm{x})=\operatorname*{\arg\min}_{\begin{subarray}{c}\hat{Y}\in 2^{\mathcal{Y}}:R_{\mathcal{T}}(\hat{Y})\leq r,\\ \hat{\omega}(y;\,\bm{x})\subseteq\hat{Y}\end{subarray}}|\hat{Y}|-\hat{P}(\hat{Y}\,|\,\bm{x})\,.$ (9)
We will refer to this optimization problem as finding the set of common ancestors. It can be thought of as a generalization of the well-known lowest common ancestor problem (Aho et al., 1973; Harel & Tarjan, 1984; Bender & Farach-Colton, 2000; Dash et al., 2013), i.e. instead of considering a single common ancestor of a set of leaves in a tree, we are looking for a set of $r$ non-overlapping ancestors whose descendants form a set of representation complexity $r$. The second term in the objective function is usually not considered in the lowest common ancestor problem, but it is needed here to make sure that the solution of (9) is unique. Since probabilities are bounded between zero and one, the second term only plays a role for sets that have the same cardinality. As an example, consider the hierarchy depicted in Figure 1 above. For $y=2$ we find $\hat{\omega}(y;\bm{x})=\{1,2,5\}$. The lowest common ancestor of this set is $A_{1}(y;\bm{x})=\{1,2,3,4,5,6,7,8\}$. The lowest set of common ancestors with representation complexity two is $A_{2}(y;\bm{x})=\{1,2,5\}$.
Let $\hat{Y}^{(1)},\ldots,\hat{Y}^{(K)}$ denote the ordered sequence of unique lowest common ancestors $A_{r}(y^{(1)};\bm{x}),\ldots,A_{r}(y^{(K)};\bm{x})$, with $y^{(1)},\ldots,y^{(K)}$ the classes sorted in decreasing order of probability for an instance $\bm{x}$. We use the ordered sequence of unique lowest common ancestors to introduce the following set-valued predictor:
$\hat{Y}_{\leq r}(\bm{x},u,\tau)=\operatorname*{\arg\max}_{\hat{Y}^{(k)}\in\{\hat{Y}^{(1)},\ldots,\hat{Y}^{(K)}\}}\big{\{}|\hat{Y}^{(k)}|:$
$\hat{P}(\hat{Y}^{(k-1)}\,|\,\bm{x})+u\cdot\hat{P}(\hat{Y}^{(k)}\setminus\hat{Y}^{(k-1)}\,|\,\bm{x})\leq\tau\big{\}}\,.$ (10)
Before discussing the set-valued predictor in (3.2), we first discuss the algorithmic aspects of computing the set of lowest common ancestors in (9), because the algorithmic insights are needed to formally prove that the sets returned by (3.2) are nested. At first impression, the combinatorial optimization problem in (9) looks challenging, but it can be solved in an efficient manner with a divide-and-conquer algorithm that is described in Algorithm S6. To this end, some data structures need to be initialized first in Algorithm S5. Both algorithms can be found in the supplementary materials.
For every internal node of the hierarchy, the optimization problem can be broken down into simpler subproblems of the same type. As an example, consider an internal node $v^{\prime}$ that has four children $(v^{\prime}_{1},v^{\prime}_{2},v^{\prime}_{3},v^{\prime}_{4})$ in a predefined hierarchy, and assume that we aim to find set of lowest common ancestors with representation complexity three for a set of classes $\hat{\omega}(y;\bm{x})$ that are all descendants of $v^{\prime}$. As potential divisions of the representation complexity, one could find one common ancestor as descendants of children $v^{\prime}_{1}$, $v^{\prime}_{3}$, and $v^{\prime}_{4}$, or as descendants of children $v^{\prime}_{1}$, $v^{\prime}_{2}$, and $v^{\prime}_{4}$, or as two descendants of $v^{\prime}_{1}$ combined with one descendant of $v^{\prime}_{4}$, etc. In fact, many combinations are possible. Speaking in formal combinatorial terminology, one has to consider for the node $v$ all weak compositions of the integer three into four elements. For each of these weak compositions, the optimization problem can be recursively divided into smaller problems with lower representation complexity, where the node $v$ is replaced by each of its non-zero children in the composition. However, implementing such a strategy in a recursive manner would heavily blow up the computations, since each of the smaller subproblems would need to be solved many times.
A dynamic programming implementation, which avoids recursion by solving the smaller subproblems in a bottom-up manner, before tackling the more challenging optimization problems higher up in the hierarchy, is able to solve (9) in an efficient manner for small values of $r$. The pseudocode of such an implementation is described in Algorithm S5 and S6. For every internal node, $r$ local optimization problems are solved by varying the local representation complexity $r_{i}$ from one to $r$, and the solutions of these optimization problems are stored. By visiting children before their parents get analysed, one can guarantee that all needed quantities are known when the different weak compositions in an internal node need to be investigated. The most critical step in the algorithm is line fifteen, i.e. where all compositions of the current representation complexity $r_{i}$ into $|T|$ elements are considered. Strictly speaking, this step has a runtime that is exponential in $r$. However, in practical situations, one would only be interested in sets with a representation complexity lower than five, so Algorithm S6 in general computes the exact solution to (9) in an efficient manner.
A similar reasoning can be made to prove that the sets returned by (3.2) are nested.
Theorem 3.1 (Nestedness of CRSVP-$r$).
Consider two threshold $\tau_{1}$ and $\tau_{2}$, with $\tau_{1}\leq\tau_{2}$, then $\hat{Y}_{\leq r}(\bm{x},u,\tau_{1})\subseteq\hat{Y}_{\leq r}(\bm{x},u,\tau_{2})$.
Proof.
Changing the threshold from $\tau_{1}$ to $\tau_{2}$ can have three consequences: (1) no additional class label can be added to $\hat{Y}_{\leq r}(\bm{x},u,\tau_{1})$, (2) the next class label from the list with sorted class probabilities is added to $\hat{Y}_{\leq r}(\bm{x},u,\tau_{1})$, or (3) more than one class label from the list with sorted class probabilities is added to $\hat{Y}_{\leq r}(\bm{x},u,\tau_{1})$. We only have to further investigate the second situation: the set-valued prediction does not change in the first situation, whereas the third situation can be reduced to the second situation by considering intermediate changes in threshold, so that only a single class is added at the same time.
For the second situation, assume that the maximally allowed representation complexity is $r$ and recall the divide-and-conquer strategy of Algorithm S6. We have to consider three subcases.
Case 1: Adding the new class $y^{\prime}$ does not change the solution of (9) found for $\tau_{1}$. This case does not need further attention, because this situation cannot occur. With the nonconformity score that is considered, the new class would have been added already with the threshold $s_{1}$, leading to a contradiction.
Case 2: Adding the new class $y^{\prime}$ changes the solution of (9) found for $\tau_{1}$, but that solution had a representation complexity that was strictly lower than the maximally allowed representation complexity $r$. The solution to the new optimization problem is the solution to the old optimization, appended with the leaf node $y^{\prime}$.
Case 3: Adding the new class $y^{\prime}$ changes the solution of (9) found for $\tau_{1}$, and that solution had a representation complexity that was equal to the maximally allowed representation complexity $r$. Adding the leaf node $y^{\prime}$ to the solution obtained for $\tau_{1}$ is in this case not possible, since this would lead to a violation of the representation complexity constraint. The new solution will be different. Let’s think about what changes are possible compared to the solution for $\tau_{1}$. When considering the dynamic programming strategy of Algorithm S6, one can observe that only a different decision can be taken in internal nodes on the path from the leaf $y^{\prime}$ to the root. Moreover, a different decision will be taken in at least one internal node on that path. In contrast, no new decision can be taken on internal nodes that are not on the path from $y^{\prime}$ to the root, because the information in these nodes has not changed. So, let’s analyze what changes are possible in the internal nodes on the path from $y^{\prime}$ to the root. In such an internal node, it can happen that the optimal weak composition for $\tau_{1}$ will change, but it can only change in such a way that the branch leading to $y^{\prime}$ is increased with one unit. As a result, one of the other branches has to decrease with one unit. Two descendants of that branch that were part of the optimal solution will have to be merged, resulting in a bigger set that corresponds to the ancestor of these nodes. Likewise, in the branch that contains $y^{\prime}$, the set size will also increase, because the new leaf $y^{\prime}$ needs to incorporated. Thus, for every internal node along the path from $y^{\prime}$ to the root, it happens that the optimal sets are nested when increasing the threshold. ∎
The pseudocode for calculating the threshold in (2.3) and set-valued predictions in (3.2) are presented in Algorithm 3 and 4. Given a hierarchical classifier and a binary tree, the computational complexity during test time is given as follows. Initialization in Algorithm S5 is given by $\mathcal{O}(K\log K)$, since in the worst case, each node in the tree must be traversed. Furthermore, calculating the set of lowest common ancestors in Algorithm S6 is given by $\mathcal{O}(K2^{r}r)$. Note that in practice, $r$ is small and the computational complexity is therefore manageable. The overall computational complexity of Algorithm 4 is then given by $\mathcal{O}(lK\log K+lK2^{r}r)$, where $l$ is the number of unique lowest common ancestors.
Algorithm 3 CRSVP-$r$ calibration – Input: $\{(\bm{x}_{i},y_{i},u_{i})\}_{i=1}^{N},r,\hat{P},\mathcal{V}_{\mathcal{T}}$, Output: Threshold in (2.3).
@@ 3.2 Conformal set-valued prediction with representation complexity (CRSVP-$r$)
3:    $\hat{Y}^{(k-1)}\leftarrow\emptyset,\hat{Y}^{(k)}\leftarrow\emptyset$
4:    $\hat{p}_{\hat{Y}^{(k-1)}}\leftarrow 0,\hat{p}_{\hat{Y}^{(k)}}\leftarrow 0$
5:    while $k\leq K$ do
6:       if $y_{i}^{(k)}\notin\hat{Y}^{(k-1)}$ then
7:          $\hat{Y}^{(k)^{\prime}}\leftarrow A_{r}(y^{(k)}_{i};\bm{x}_{i})$ (by means of Algorithm S5 and  S6)
@@
16:       $k\leftarrow k+1$
17:    end while
18:    $\tau_{i}\leftarrow\hat{p}_{\hat{Y}^{(k)}}-u_{i}\cdot(\hat{p}_{\hat{Y}^{(k)}}-\hat{p}_{\hat{Y}^{(k-1)}})$
19:  end for
20:  $\tau^{*}\leftarrow$ the $\lceil(1-\alpha)(N+1)\rceil$-th largest value in $\{\tau_{i}\}_{i=1}^{N}$
21:  return $\tau^{*}$
Algorithm 4 CRSVP-$r$ inference – Input: $\bm{x},\tau^{*},u,r,\hat{P},\mathcal{V}_{\mathcal{T}}$, Output: Set-valued prediction in (3.2).
1:  $k\leftarrow 1$
2:  $\hat{Y}^{(k-1)}\leftarrow\emptyset,\hat{Y}^{(k)}\leftarrow\emptyset$
3:  $\hat{p}_{\hat{Y}^{(k-1)}}\leftarrow 0,\hat{p}_{\hat{Y}^{(k)}}\leftarrow 0$
4:  while $k\leq K$ do
@@
6:       $\hat{Y}^{(k)^{\prime}}\leftarrow A_{r}(y^{(k)};\bm{x})$ (by means of Algorithm S5 and  S6)
@@
17:  return $\hat{Y}^{(k-1)}$

### v3

@@ 3 Proposed methods
Building upon the concepts in Section 2.3, we are now ready to discuss two algorithms that provide valid set-valued predictions for hierarchical classification that satisfy (1) above. The first algorithm provides marginal validity guarantees in classical hierarchical classification settings, i.e. where sets are restricted to nodes of the tree structure, thus $R_{\mathcal{T}}(\hat{Y})=1$. The second algorithm provides marginal validity guarantees in restricted set-valued prediction settings, i.e. where the representation complexity of the predicted set is restricted to $R_{\mathcal{T}}(\hat{Y})\leq r$ for a user-defined value $r$. Note that traditional conformal prediction in flat classification is obtained for the second algorithm when there is no restriction on the representation complexity (e.g. $r=K$).
3.1 Conformal restricted set-valued prediction
Algorithm 1 CRSVP calibration – Input: $\{(\bm{x}_{i},y_{i},u_{i})\}_{i=1}^{N},\hat{P},\mathcal{V}_{\mathcal{T}}$, Output: Threshold in (5).
@@ 3.1 Conformal restricted set-valued prediction
5:   $\hat{p}_{\hat{Y}^{\prime}}\leftarrow\hat{p}_{\hat{Y}}$, $\hat{p}_{\hat{Y}}\leftarrow\hat{P}(\mathrm{pa}(\hat{Y})\,|\,\bm{x}_{i})$
[…]
@@
[…]
@@
Algorithm 2 CRSVP inference – Input: $\bm{x},\tau^{*},u,\hat{P},\mathcal{V}_{\mathcal{T}}$, Output: Set-valued prediction in (7).
[…]
3:while $\mathrm{pa}(\hat{Y}^{\prime})\neq\emptyset$ do
4:  if $\hat{p}_{\hat{Y}}\geq\tau^{*}$ then
@@
7:  $\hat{p}_{\hat{Y}^{\prime}}\leftarrow\hat{p}_{\hat{Y}}$, $\hat{p}_{\hat{Y}}\leftarrow\hat{P}(\mathrm{pa}(\hat{Y})\,|\,\bm{x})$
[…]
@@
10:if $\hat{p}_{\hat{Y}}-u\cdot(\hat{p}_{\hat{Y}}-\hat{p}_{\hat{Y}^{\prime}})\geq\tau^{*}$ then
11:  return $\hat{Y}^{\prime}$
12:else
13:  return $\hat{Y}$
14:end if
The first approach, called conformal restricted set-valued prediction (CRSVP), predicts sets that are restricted to nodes of the hierarchy, thus having representation complexity $R_{\mathcal{T}}(\hat{Y})=1$. Assume one has fitted a classifier using a training dataset as described in Section 2.1, and let $\hat{y}(\bm{x})$ denote the mode of the distribution $\hat{P}(.\,|\,\bm{x})$, i.e. the leaf node with highest probability mass. For our first approach, we consider the following set-valued predictor:
$\hat{Y}_{1}(\bm{x},u,\tau)=\operatorname*{\arg\max}_{\hat{Y}\in\mathrm{Path}(\hat{y}(\bm{x}))}\big\{|\hat{Y}|:\hat{P}(\hat{Y}\,|\,\bm{x})+\\ u\cdot\hat{P}(\mathrm{pa}(\hat{Y})\setminus\hat{Y}\,|\,\bm{x})\leq\tau\big\}\,$, (7)
with $\mathrm{Path}(v)$ the path to the root, and $\mathrm{pa}(v)$ the parent of a node $v$, as defined in Section 2. $\hat{P}(\hat{Y}\,|\,\bm{x})$ can be computed using (2), but also observe that the probability mass of an internal node $v$ corresponds to $\hat{P}(v\,|\,\bm{x})=\sum_{c\in v}\hat{P}(c\,|\,\bm{x})$, i.e. the sum of the probability masses of the leaf nodes that are descendants of $v$. However, computing the probability mass of an internal node and its associated set of classes in this way is computationally less efficient than using the chain rule in (2), when hierarchical classifiers have been fitted during training.
It is clear that the set-valued predictor in (7) satisfies the nestedness property in (4). Indeed, starting from the mode of the distribution $\hat{y}(\bm{x})$, every node on the path connecting that node and the root is nested by definition of the hierarchical tree structure $\mathcal{T}$. In line with Angelopoulos et al., (2020), the first term increases as we move up the tree, while the second term contains a random draw from the uniform distribution $\mathcal{U}(0,1)$, for handling discrete jumps in probability mass when following the path towards the root. The randomization is needed to prevent over-coverage.
The algorithm for calculating the threshold in (5) using the calibration dataset is presented in Algorithm 1. This algorithm first computes the mode of the estimated class probabilities for each instance in the calibration dataset. For small hierarchies, this can be done using exhaustive search, but more efficient inference algorithms have been developed when the computational cost becomes a burden (Dembczyński et al.,, 2012; Kumar et al.,, 2013; Mena Waldo et al.,, 2015). Subsequently, starting from the mode, Algorithm 1 computes for every calibration instance the internal node, on the path to the root, that also includes the true class label. Nonconformity scores that incorporate a randomization component are constructed for these internal nodes, and the final threshold is deduced from the scores. Algorithm 1 has a worst case $\mathcal{O}(N\,K)$ time complexity when exhaustive search is applied in the first step.
The pseudocode for computing the set-valued prediction for a new test instance, as defined in (7), is shown in Algorithm 2. Given a hierarchical classifier, the computational complexity during test time is given by $\mathcal{O}(\log K)$.
3.2 Conformal set-valued prediction with representation complexity
Algorithm 3 CRSVP-$r$ calibration – Input: $\{(\bm{x}_{i},y_{i},u_{i})\}_{i=1}^{N},r,\hat{P},\mathcal{V}_{\mathcal{T}}$, Output: Threshold in (5).
@@ 3.2 Conformal set-valued prediction with representation complexity
3:  $\hat{Y}^{(k-1)}\leftarrow\emptyset,\hat{Y}^{(k)}\leftarrow\operatorname*{\arg\max}_{c\in\mathcal{Y}}\hat{P}(c\,|\,\bm{x}_{i})$
4:  $\hat{p}_{\hat{Y}^{(k-1)}}\leftarrow 0,\hat{p}_{\hat{Y}^{(k)}}\leftarrow\hat{P}(\hat{Y}^{(k)}\,|\,\bm{x}_{i})$
5:  while $k<K$ do
6:   if $y_{i}\notin\hat{Y}^{(k)}$ then
7:    $\hat{Y}^{(k)^{\prime}}\leftarrow A_{r}(S_{k};\bm{x}_{i})$ (by means of Algorithm 5)
@@
16:  end while
17:  $\tau_{i}\leftarrow\hat{p}_{\hat{Y}^{(k)}}-u_{i}\cdot(\hat{p}_{\hat{Y}^{(k)}}-\hat{p}_{\hat{Y}^{(k-1)}})$
18:end for
[…]
20:return $\tau^{*}$
Algorithm 4 CRSVP-$r$ inference – Input: $\bm{x},\tau^{*},u,r,\hat{P},\mathcal{V}_{\mathcal{T}}$, Output: Set-valued prediction in (9).
$1k\leftarrow$ 1
[…]
4:while $k<K$ do
@@
6:   $\hat{Y}^{(k)^{\prime}}\leftarrow A_{r}(S_{k};\bm{x})$ (by means of Algorithm 5)
@@
17:if $\hat{p}_{\hat{Y}^{(k)}}-u\cdot(\hat{p}_{\hat{Y}^{(k)}}-\hat{p}_{\hat{Y}^{(k-1)}})\geq\tau^{*}$ then
18:  return $\hat{Y}^{(k-1)}$
19:else
20:  return $\hat{Y}^{(k)}$
21:end if
Algorithm 5 Dynamic programming solution for set of lowest common ancestors – Input: $\bm{x},S_{k},r,\hat{P},\mathcal{V}_{\mathcal{T}},A_{r}(S_{k-1};\bm{x})$, Output: $A_{r}(S_{k};\bm{x})$
[…]
2:for all $v\in\mathcal{V}_{\mathcal{T}}$ do
3:  $s^{1}(v)\leftarrow\emptyset,\ldots,s^{r}(v)\leftarrow\emptyset$
4:  if $|v|=1\;\wedge\;v\cap S_{k}\neq\emptyset$ then
5:   if $\mathrm{pa}(v)\notin\mathcal{Q}$ then
6:    $\mathcal{Q}\mathrm{.add}(\mathrm{pa}(v))$
7:   end if
8:   $s^{1}(v)\leftarrow v,\ldots,s^{r}(v)\leftarrow v$ $\triangleright$ initialize leaves that overlap with $S_{k}$
9:  end if
10:end for
11:while $\mathcal{Q}$ is not empty do
12:  $v\leftarrow\mathcal{Q}$.pop()
13:  $T\leftarrow\{v^{\prime}:v^{\prime}\in\mathrm{ch}(v)\wedge v^{\prime}\cap S_{k}\neq\emptyset\}$ $\triangleright$ visit all children that overlap with $S_{k}$
14:  for $r_{i}=1$ to $r$ do
15:   if $|T|>r_{i}$ then
16:    $s^{r_{i}}(v)\leftarrow v$
17:   else
18:    $M\leftarrow$ all compositions of $r_{i}$ into $|T|$ elements
19:    $u^{*}\leftarrow+\infty$
20:    for $(i_{0},\ldots,i_{|T|-1})\in M$ do
21:     $s\leftarrow s^{i_{0}}(T[0])\cup\cdots\cup s^{i_{|T|-1}}(T[|T|-1])$
22:     if $|s|-\hat{p}_{s}<u^{*}$ then
23:      if $v$ is the root then
24:       if $A_{r}(S_{k-1};\bm{x})\subseteq s$ then
25:        $s^{r_{i}}(v)\leftarrow s$
26:        $u^{*}\leftarrow|s|-\hat{p}_{s}$
27:       end if
28:      else
29:       $s^{r_{i}}(v)\leftarrow s$
30:       $u^{*}\leftarrow|s|-\hat{p}_{s}$
31:      end if
32:     end if
33:    end for
34:   end if
35:  end for
36:  if $\mathrm{pa}(v)\notin\mathcal{Q}\wedge\mathrm{pa}(v)\neq\emptyset$ then
37:   $\mathcal{Q}\mathrm{.add}(\mathrm{pa}(v))$
38:  end if
39:end while
40:return $s^{r}(v_{1})$
It should be obvious that sets of representation complexity one quickly become very big, since they correspond to internal nodes of the hierarchy that exhibit a predefined coverage level. Therefore, we present a second approach, called conformal set-valued prediction with representation complexity (CRSVP-$r$), that relaxes the representation complexity constraint to $R_{\mathcal{T}}(\hat{Y})\leq r$ with $r$ a user-defined parameter. Let $S_{k}\equiv\{y^{(1)},\ldots,y^{(k)}\}$ denote the top-$k$ classes for $\bm{x}$ when sorting the classes according to $\hat{P}(y\,|\,\bm{x})$, similar as in (6). We define the following sequence of optimization problems for any $k\in\{1,...,K\}$:
$A_{r}(S_{k};\bm{x})=\operatorname*{\arg\min}_{\begin{subarray}{c}\hat{Y}\in 2^{\mathcal{Y}}:R_{\mathcal{T}}(\hat{Y})\leq r,\\ A_{r}(S_{k-1};\bm{x})\,\cup\,y^{(k)}\subseteq\hat{Y}\end{subarray}}|\hat{Y}|-\hat{P}(\hat{Y}\,|\,\bm{x})\,$, (8)
with $A_{r}(S_{1};\bm{x})=y^{(1)}$. Remark that these optimization problems have to be solved in the correct order, because of the nestedness condition. For fixed $k$ we refer to this optimization problem as finding the set of common ancestors. It can be thought of as a variant of the well-known lowest common ancestor problem (Aho et al.,, 1973; Harel and Tarjan,, 1984; Bender and Farach-Colton,, 2000; Dash et al.,, 2013), i.e. instead of considering a single common ancestor of a set of leaves in a tree, we are looking for a set of $r$ non-overlapping ancestors whose descendants form a set of representation complexity $r$. Moreover, we introduce the nested set condition, and the second term in the objective function is also not considered in the lowest common ancestor problem. Those two changes are needed to make sure that (1) the solutions for increasing $k$ are nested, and (2) the solution for fixed $k$ is unique. Since probabilities are bounded between zero and one, the second term only plays a role for sets that have the same cardinality (i.e. the two terms are lexicographically ordered). As an example, consider the hierarchy depicted in Figure 2. For $S_{3}=\{1,2,5\}$, the lowest common ancestor is given by $A_{1}(S_{3};\bm{x})=\{1,2,3,4,5,6,7,8\}$. The lowest set of common ancestors with representation complexity two is $A_{2}(S_{3};\bm{x})=\{1,2,5\}$, illustrating that set size can be drastically reduced by increasing the representation complexity.
Let $\hat{Y}^{(1)},\ldots,\hat{Y}^{(K)}$ denote the ordered sequence of unique lowest common ancestors $A_{r}(S_{1};\bm{x}),\ldots,A_{r}(S_{K};\bm{x})$. We use this sequence to introduce the following set-valued predictor:
$\hat{Y}_{\leq r}(\bm{x},u,\tau)=\operatorname*{\arg\max}_{\hat{Y}^{(k)}\in\{\hat{Y}^{(1)},\ldots,\hat{Y}^{(K)}\}}\big\{|\hat{Y}^{(k)}|:\\ \hat{P}(\hat{Y}^{(k-1)}\,|\,\bm{x})+u\cdot\hat{P}(\hat{Y}^{(k)}\setminus\hat{Y}^{(k-1)}\,|\,\bm{x})\leq\tau\big\}\,.$ (9)
Due to the nested set constraint in (8), the set-valued predictor in (9) satisfies the nestedness property in (4). Thus, this set-valued predictor automatically has the classical marginal coverage guarantees of conformal prediction. Similar as for the first approach, the randomization term $u\cdot\hat{P}(\hat{Y}^{(k)}\setminus\hat{Y}^{(k-1)}\,|\,\bm{x})$ is again included to prevent over-coverage.
Let us now discuss the algorithmic aspects of computing the set of lowest common ancestors in (8). At first impression, the combinatorial optimization problem in (8) looks challenging, but it can be solved in an efficient manner with a divide-and-conquer algorithm that is described in Algorithm 5.
For every internal node of the hierarchy, the optimization problem can be broken down into simpler subproblems of the same type. As an example, consider in a predefined hierarchy an internal node $v^{\prime}$ that has four children $(v^{\prime}_{1},v^{\prime}_{2},v^{\prime}_{3},v^{\prime}_{4})$ containing at least one element of $S_{k}$, and assume that we aim to find the set of lowest common ancestors with representation complexity three for a set of classes $S_{k}$ that are all descendants of $v^{\prime}$. As potential divisions of the representation complexity, one could find one common ancestor as descendants of children $v^{\prime}_{1}$, $v^{\prime}_{3}$, and $v^{\prime}_{4}$, or as descendants of children $v^{\prime}_{1}$, $v^{\prime}_{2}$, and $v^{\prime}_{4}$, or as two descendants of $v^{\prime}_{1}$ combined with one descendant of $v^{\prime}_{4}$, etc. In fact, many combinations are possible. Speaking in formal combinatorial terminology, one has to consider for the node $v$ all compositions of the integer three into four elements. For each of these compositions, the optimization problem can be recursively divided into smaller problems with lower representation complexity, where the node $v$ is replaced by each of its non-zero children in the composition. However, implementing such a strategy in a recursive manner would heavily blow up the computations, since each of the smaller subproblems would need to be solved many times.
A dynamic programming implementation, which avoids recursion by solving the smaller subproblems in a bottom-up manner, before tackling the more challenging optimization problems higher up in the hierarchy, is able to solve (8) in an efficient manner for small values of $r$. The pseudocode of such an implementation is described in Algorithm 5. For every internal node, $r$ local optimization problems are solved by varying the local representation complexity $r_{i}$ from one to $r$, and the solutions of these optimization problems are stored. By visiting children before their parents get analysed, one can guarantee that all needed quantities are known when the different compositions in an internal node need to be investigated. The most critical step in Algorithm 5 is line 20, where all compositions of the current representation complexity $r_{i}$ into $|T|$ elements are considered (with $T$ the set of all children that contain at least one element of $S_{k}$). Strictly speaking, this step has a runtime that is exponential in $r$. However, in practical situations, one would only be interested in sets with a representation complexity lower than five, so Algorithm 5 in general computes the exact solution to (8) in an efficient manner.
The pseudocode for calculating the threshold in (5) and set-valued predictions in (9) is presented in Algorithm 3 and 4. The computational complexity during test time is dominated by the computation of the set of lowest common ancestors in (8), which needs to be computed for every $k$ until the stopping criterion in Algorithm 4 is satisfied. Furthermore, the worst-case time complexity of calculating the set of lowest common ancestors in Algorithm 5 is upper bounded by $\mathcal{O}(K^{2}r^{d})$, with $d$ the maximum out-degree of the tree $\mathcal{T}$. This is small for regimes that are practically useful (fixed $r\leq 3$ and moderate $d$).